# RS-Agent: Automating Remote Sensing Tasks through Intelligent Agent

### v2

@@ 4.1 Implementation Details
To ensure the flexibility and generalizability of RS-Agent, we build the Central Controller with LangChain framework [Chase, 2022] and general-purpose large language models. Unless otherwise specified, RS-Agent uses the Qwen 2.5 32B-Instruct model [Yang et al., 2024], balancing the model capability with inference speed.
Within the Task-Aware Retrieval method, we employ the open-source m3e-base [Chen et al., 2024] model as the embedding model to convert textual data into vectors. These embeddings are then indexed using the FAISS [Jégou et al., 2022] algorithm, enabling efficient retrieval of relevant information. For DualRAG method, we return the top-$N$ most relevant knowledge documents, with $N$=30.
@@ 4.2.1 Task Planning Accuracy Compared to SOTA Agent
Dataset. We construct a dedicated dataset to evaluate RS-Agent’s task planning capabilities, measuring accuracy by whether the first tool invoked matches the correct one for the query. To ensure a fair comparison with RS-ChatGPT [Guo et al., 2024a], we adopt seven core tasks from RS-ChatGPT: scene classification, land use segmentation, object detection, image captioning, edge detection, object counting, and instance segmentation. For each task, we create average 20 query-solution pairs, including both clear and intentionally ambiguous instructions to assess the model’s ability to interpret user intent, resolve ambiguity, and select the appropriate tool.
@@ 4.3.1 Object Counting
Datasets. We use the validation set from DOTAv1 [Xia et al., 2018] to assess the object counting ability. The set contains 458 images and 1,099 questions. For the object counting task, we used the prompt: “What is the number of the object? Answer the question using a number.” The object could be any one of 15 typical remote sensing target types, such as plane, ship and bridge.
Metric. We evaluate object-counting performance using three metrics: absolute accuracy, interval-matching accuracy, and relative error. Absolute accuracy measures the percentage of predictions that exactly match the ground truth. Interval-matching accuracy considers a prediction correct if it falls within the same predefined range as the ground truth. The intervals are: 0, 1–10, 11–100, 101–1000, and $>$ 1000. Relative error $e_{r}$ quantifies the deviation of the prediction from the ground truth and is computed as follows:
@@
Results. As presented in Table 3, RS-Agent outperforms both GeoChat [Kuckreja et al., 2023] and LHRS-Bot [Muhtar et al., 2024] across all evaluation metrics. It achieves an absolute accuracy of 33.30%, significantly surpassing GeoChat (17.65%) and LHRS-Bot (14.92%). Its interval-matching accuracy reaches 75.98%, slightly higher than GeoChat (74.61%) and notably better than LHRS-Bot (60.78%). In terms of relative error, RS-Agent achieves a score of 0.28—substantially lower than GeoChat (0.65) and LHRS-Bot (0.56)—indicating more precise and consistent predictions. These results demonstrate the robustness and reliability of RS-Agent in object counting tasks.
@@
LLaVA-1.5 [Liu et al., 2023] 59.22% 73.16% 65.19% 65.86%
MiniGPTv2 [Chen et al., 2023] 60.02% 51.64% 67.64% 59.77%
InstructBLIP [Dai et al., 2024] 62.62% 48.83% 63.92% 59.12%
mPLUG-Owl2 [Ye et al., 2023] 57.99% 74.04% 65.04% 65.69%
QWen-VL-Chat [Bai et al., 2023] 62.00% 47.65% 54.64% 58.73%
RSVQA [Lobry et al., 2020] 90.00% 87.47% 81.50% 86.32%
SkyEyeGPT [Zhan et al., 2025] 88.93% 88.63% 75.00% 84.16%
LHRS-Bot [Muhtar et al., 2024] 89.07% 88.51% 90.00% 89.19%
GeoChat [Kuckreja et al., 2023] 94.00% 91.09% 90.33% 90.70%
@@
MiniGPTv2 [Chen et al., 2023] - 52.60% 62.90%
Qwen-VL-Chat [Bai et al., 2023] - 52.60% 62.90%
LLaVA-1.5 [Liu et al., 2023] - 51.00% 68.00%
GeoChat [Kuckreja et al., 2023] 51.43% 72.03% 84.43%
LHRS-Bot [Muhtar et al., 2024] 63.85% 91.29% 96.63%
@@ 4.3.2 Visual Question Answering
Datasets. We evaluate VQA performance using the widely adopted RSVQA-LR dataset [Lobry et al., 2020], which comprises image–question–answer triplets tailored for remote sensing visual question answering. The dataset includes 772 satellite images and 77,232 question–answer pairs, covering four question types: presence, comparison, rural/urban classification, and counting.
Compared Methods. We compare RS-Agent with two groups of multimodal large language models: (1) general-purpose models trained on broad, non-specialized corpora, e.g., LLaVA-1.5 [Liu et al., 2023], MiniGPTv2 [Chen et al., 2023], InstructBLIP [Dai et al., 2024], mPLUG-Owl2 [Ye et al., 2023] and QWen-VL-Chat [Bai et al., 2023] and (2) remote sensing-specific models fine-tuned on domain-specific remote sensing datasets, e.g., RSVQA [Lobry et al., 2020], SkyEyeGPT [Zhan et al., 2025], LHRS-Bot [Muhtar et al., 2024] and GeoChat [Kuckreja et al., 2023].
@@ 4.3.3 Scene Classification
Datasets. For evaluation, we utilized the RSSDIVCS [Li et al., 2021b] dataset, AID [Xia et al., 2017] dataset, and UCMerced [Yang and Newsam, 2010] dataset. The AID comprises a substantial collection of aerial images sourced from Google Earth, spanning 30 classes. The UCMerced consists of 2100 images across 21 land use scene categories.
To perform this task, RS-Agent is prompted with: “Identify the scene depicted in this image.” For GeoChat [Kuckreja et al., 2023] and LHRS-Bot [Muhtar et al., 2024], the prompt used is: “Classify the image within one of the given classes: dense residential area, …, school. Answer with one word or short phrase.”
For scene classification, RS-Agent employs a Vision Transformer (ViT-B16)[Dosovitskiy et al., 2020] pretrained with self-supervision on the unlabelled ImageNet dataset[Deng et al., 2009] as the backbone. The model is then fine-tuned on remote sensing imagery to function as the scene classifier. Training is conducted on the RSSDIVCS dataset [Li et al., 2021b], which comprises 70 scene categories, each containing 800 images at a resolution of 256$\times$256 pixels. We use 80% of the images from each category for training.
Results. As shown in Table 5, RS-Agent consistently outperforms baseline models across all datasets. On the RSSDIVCS [Li et al., 2021b] dataset, it achieves 98.00% accuracy, significantly higher than GeoChat (51.43%) and LHRS-Bot (63.85%). On AID [Xia et al., 2017], RS-Agent reaches 96.88%, surpassing GeoChat (72.03%) and LHRS-Bot (91.29%). Similarly, on the UCMerced [Yang and Newsam, 2010] dataset, RS-Agent attains 98.63%, outperforming GeoChat (84.43%) and LHRS-Bot (96.63%).
@@
Baseline+$T_{\text{inf}}$ 52.63% 36.84% 95.24% 95.24% 95.24% 100% 100% 82.01% 84.34%
Baseline+$S_{\text{ret}}$ 52.63% 31.58% 100% 80.95% 52.38% 94.74% 94.74% 73.38% 84.57%
Baseline+$T_{\text{inf}}$+$S_{\text{ret}}$ 100% 84.21% 90.48% 95.24% 100% 84.21% 100% 93.53% 90.88%
@@ 4.4 Effectiveness of the Task-Aware Retrieval method
  • Baseline+$T_{\text{inf}}$: The RS-Agent with Solution Space, where we only perform the Task Inference.
  • Baseline+$S_{\text{retr}}$: The RS-Agent with Solution Space, where we only retrieves task-specific guidance by directly using $Q$ as the retrieval input, allowing LLM to leverage external knowledge when interpreting the task.
  • Baseline+$T_{\text{inf}}$+$S_{\text{retr}}$: Our full RS-Agent network.
@@
cloud_removal [Pan, 2020] Cloud removal from satellite images Remove the clouds in this image.
image_dehazing [Luo et al., 2023] Haze removal from images Dehaze this foggy image.
super_resolution [Wang et al., 2021] Image super-resolution (2×) Enhance the resolution of this image.
denoising [Vaksman et al., 2020] Image denoising Remove noise from this image.
caption [Kuckreja et al., 2023] Geo-specific VQA and captioning What is in this remote sensing image?
optical_detection [Ultralytics, 2023] Optical image target detection Detect objects in this optical image.
@@
sar_detection [Zhou et al., 2024] Target detection in SAR images Find the objects in this SAR image.
@@
building_damage_detection [Zheng et al., 2021] Building damage assessment Which buildings are damaged?
building_extraction [He et al., 2023] Building extraction from images Extract all buildings from the image.
road_extraction [Hetang et al., 2024] Road extraction from images Extract roads from the scene.
horizontal_object_detection [Wang et al., 2024b] Horizontal bounding box detection Detect objects using horizontal boxes.
rotated_object_detection [Wang et al., 2024b] Rotated object detection Detect objects using rotated boxes.
semantic_segmentation [Wang et al., 2024b] Pixel-wise semantic segmentation Segment the different regions in this image.
land_use_classification [Wang et al., 2022] Land use categorization What are the land use types in this image?
@@ 4.5 Effectiveness of the DualRAG method
In this Section, to verify the effectiveness of DualRAG method, we perform an in-depth comparision between RS-Agent with DualRAG and LightRAG [Guo et al., 2024b].
Datasets. In this study, we utilize two datasets: the RSaircraft dataset, which we have specifically constructed for remote sensing applications, and the Mix dataset from the UltraDomain benchmark [Qian et al., 2024], a general-domain RAG dataset. The RSaircraft dataset focuses on aircraft-related information, tailored for remote sensing tasks, and was built by systematically collecting and processing structured data from publicly available sources. The Mix dataset, on the other hand, consists of a diverse collection of literary, biographical, and philosophical texts, representing a broad range of disciplines such as cultural studies, historical analysis, and philosophical discourse. The inclusion of both datasets allows us to evaluate the performance of DualRAG in both domain-specific (remote sensing) and general-domain contexts.
Metric. Following the approach outlined in LightRAG [Guo et al., 2024b], we employ GPT-4o-mini as an automated evaluator to assess model responses across four dimensions: Comprehensiveness, Diversity, Empowerment, and Overall quality. The evaluation is carried out in a pairwise comparison setting, where the judge selects the superior response between two candidates. The usage of keywords follows three modes: Local, Global, and Hybrid, as defined in LightRAG.
Results. Table 7 presents the win rates of LightRAG [Guo et al., 2024b] compared to our RS-Agent with DualRAG across two datasets (RSaircraft and Mix) and four evaluation dimensions. RS-Agent consistently outperforms LightRAG across nearly all settings, with particularly substantial improvements observed on the RSaircraft dataset.

### v3

@@ 4.1 Implementation Details
To ensure the flexibility and generalizability of RS-Agent, we build the Central Controller with LangChain framework [5] and general-purpose large language models. Unless otherwise specified, RS-Agent uses the Qwen 2.5 32B-Instruct model [67], balancing the model capability with inference speed.
Within the Task-Aware Retrieval method, we employ the open-source m3e-base [6] model as the embedding model to convert textual data into vectors. These embeddings are then indexed using the FAISS [23] algorithm, enabling efficient retrieval of relevant information. For DualRAG method, we return the top-$N$ most relevant knowledge documents, with NN=30.
@@ 4.2.1 Task Planning Accuracy Compared to SOTA Agent
Dataset. We construct a dedicated dataset to evaluate RS-Agent’s task planning capabilities, measuring accuracy by whether the first tool invoked matches the correct one for the query. To ensure a fair comparison with RS-ChatGPT [18], we adopt seven core tasks from RS-ChatGPT: scene classification, land use segmentation, object detection, image captioning, edge detection, object counting, and instance segmentation. For each task, we create average 20 query-solution pairs, including both clear and intentionally ambiguous instructions to assess the model’s ability to interpret user intent, resolve ambiguity, and select the appropriate tool.
@@ 4.3.1 Object Counting
Datasets. We use the validation set from DOTAv1 [64] to assess the object counting ability. The set contains 458 images and 1,099 questions. For the object counting task, we used the prompt: “What is the number of the object? Answer the question using a number.” The object could be any one of 15 typical remote sensing target types, such as plane, ship and bridge.
Metric. We evaluate object-counting performance using three metrics: absolute accuracy, interval-matching accuracy, and relative error. Absolute accuracy measures the percentage of predictions that exactly match the ground truth. Interval-matching accuracy considers a prediction correct if it falls within the same predefined range as the ground truth. The intervals are: 0, 1–10, 11–100, 101–1000, and >> 1000. Relative error $e_{r}$ quantifies the deviation of the prediction from the ground truth and is computed as follows:
@@
Results. As presented in Table 3, RS-Agent outperforms both GeoChat [25] and LHRS-Bot [36] across all evaluation metrics. It achieves an absolute accuracy of 33.30%, significantly surpassing GeoChat (17.65%) and LHRS-Bot (14.92%). Its interval-matching accuracy reaches 75.98%, slightly higher than GeoChat (74.61%) and notably better than LHRS-Bot (60.78%). In terms of relative error, RS-Agent achieves a score of 0.28—substantially lower than GeoChat (0.65) and LHRS-Bot (0.56)—indicating more precise and consistent predictions. These results demonstrate the robustness and reliability of RS-Agent in object counting tasks.
@@
LLaVA-1.5 [31] 59.22% 73.16% 65.19% 65.86%
MiniGPTv2 [7] 60.02% 51.64% 67.64% 59.77%
InstructBLIP [8] 62.62% 48.83% 63.92% 59.12%
mPLUG-Owl2 [69] 57.99% 74.04% 65.04% 65.69%
QWen-VL-Chat [1] 62.00% 47.65% 54.64% 58.73%
RSVQA [32] 90.00% 87.47% 81.50% 86.32%
SkyEyeGPT [70] 88.93% 88.63% 75.00% 84.16%
LHRS-Bot [36] 89.07% 88.51% 90.00% 89.19%
GeoChat [25] 94.00% 91.09% 90.33% 90.70%
@@
MiniGPTv2 [7] - 52.60% 62.90%
Qwen-VL-Chat [1] - 52.60% 62.90%
LLaVA-1.5 [31] - 51.00% 68.00%
GeoChat [25] 51.43% 72.03% 84.43%
LHRS-Bot [36] 63.85% 91.29% 96.63%
@@ 4.3.2 Visual Question Answering
Datasets. We evaluate VQA performance using the widely adopted RSVQA-LR dataset [32], which comprises image–question–answer triplets tailored for remote sensing visual question answering. The dataset includes 772 satellite images and 77,232 question–answer pairs, covering four question types: presence, comparison, rural/urban classification, and counting.
Compared Methods. We compare RS-Agent with two groups of multimodal large language models: (1) general-purpose models trained on broad, non-specialized corpora, e.g., LLaVA-1.5 [31], MiniGPTv2 [7], InstructBLIP [8], mPLUG-Owl2 [69] and QWen-VL-Chat [1] and (2) remote sensing-specific models fine-tuned on domain-specific remote sensing datasets, e.g., RSVQA [32], SkyEyeGPT [70], LHRS-Bot [36] and GeoChat [25].
@@ 4.3.3 Scene Classification
Datasets. For evaluation, we utilized the RSSDIVCS [27] dataset, AID [65] dataset, and UCMerced [68] dataset. The AID comprises a substantial collection of aerial images sourced from Google Earth, spanning 30 classes. The UCMerced consists of 2100 images across 21 land use scene categories.
To perform this task, RS-Agent is prompted with: “Identify the scene depicted in this image.” For GeoChat [25] and LHRS-Bot [36], the prompt used is: “Classify the image within one of the given classes: dense residential area, …, school. Answer with one word or short phrase.”
For scene classification, RS-Agent employs a Vision Transformer (ViT-B16)[11] pretrained with self-supervision on the unlabelled ImageNet dataset[9] as the backbone. The model is then fine-tuned on remote sensing imagery to function as the scene classifier. Training is conducted on the RSSDIVCS dataset [27], which comprises 70 scene categories, each containing 800 images at a resolution of $\times256$ pixels. We use 80% of the images from each category for training.
Results. As shown in Table 5, RS-Agent consistently outperforms baseline models across all datasets. On the RSSDIVCS [27] dataset, it achieves 98.00% accuracy, significantly higher than GeoChat (51.43%) and LHRS-Bot (63.85%). On AID [65], RS-Agent reaches 96.88%, surpassing GeoChat (72.03%) and LHRS-Bot (91.29%). Similarly, on the UCMerced [68] dataset, RS-Agent attains 98.63%, outperforming GeoChat (84.43%) and LHRS-Bot (96.63%).
@@
[…]
@@ 4.4 Effectiveness of the Task-Aware Retrieval method
[…]
@@
cloud_removal [37] Cloud removal from satellite images Remove the clouds in this image.
image_dehazing [34] Haze removal from images Dehaze this foggy image.
super_resolution [61] Image super-resolution (2×) Enhance the resolution of this image.
denoising [52] Image denoising Remove noise from this image.
caption [25] Geo-specific VQA and captioning What is in this remote sensing image?
optical_detection [51] Optical image target detection Detect objects in this optical image.
@@
sar_detection [72] Target detection in SAR images Find the objects in this SAR image.
@@
building_damage_detection [71] Building damage assessment Which buildings are damaged?
building_extraction [20] Building extraction from images Extract all buildings from the image.
road_extraction [21] Road extraction from images Extract roads from the scene.
horizontal_object_detection [54] Horizontal bounding box detection Detect objects using horizontal boxes.
rotated_object_detection [54] Rotated object detection Detect objects using rotated boxes.
semantic_segmentation [54] Pixel-wise semantic segmentation Segment the different regions in this image.
land_use_classification [58] Land use categorization What are the land use types in this image?
@@ 4.5 Effectiveness of the DualRAG method
In this Section, to verify the effectiveness of DualRAG method, we perform an in-depth comparision between RS-Agent with DualRAG and LightRAG [19].
Datasets. In this study, we utilize two datasets: the RSaircraft dataset, which we have specifically constructed for remote sensing applications, and the Mix dataset from the UltraDomain benchmark [39], a general-domain RAG dataset. The RSaircraft dataset focuses on aircraft-related information, tailored for remote sensing tasks, and was built by systematically collecting and processing structured data from publicly available sources. The Mix dataset, on the other hand, consists of a diverse collection of literary, biographical, and philosophical texts, representing a broad range of disciplines such as cultural studies, historical analysis, and philosophical discourse. The inclusion of both datasets allows us to evaluate the performance of DualRAG in both domain-specific (remote sensing) and general-domain contexts.
Metric. Following the approach outlined in LightRAG [19], we employ GPT-4o-mini as an automated evaluator to assess model responses across four dimensions: Comprehensiveness, Diversity, Empowerment, and Overall quality. The evaluation is carried out in a pairwise comparison setting, where the judge selects the superior response between two candidates. The usage of keywords follows three modes: Local, Global, and Hybrid, as defined in LightRAG.
Results. Table 7 presents the win rates of LightRAG [19] compared to our RS-Agent with DualRAG across two datasets (RSaircraft and Mix) and four evaluation dimensions. RS-Agent consistently outperforms LightRAG across nearly all settings, with particularly substantial improvements observed on the RSaircraft dataset.